# Sparsing Law: Towards Large Language Models with Greater Activation Sparsity

### v1

@@ Abstract
Activation sparsity denotes the existence of substantial weakly-contributed elements within activation outputs that can be eliminated, benefiting many important applications concerned with large language models (LLMs), such as computation acceleration and model interpretability. Although promoting greater activation sparsity within LLMs deserves deep studies, existing works lack comprehensive and quantitative research on the correlation between activation sparsity and potentially influential factors. In this paper, we present a comprehensive study on the quantitative scaling properties and influential factors of the activation sparsity within decoder-only Transformer-based LLMs. Specifically, we propose PPL-$p\%$ sparsity, a precise and performance-aware activation sparsity metric that is applicable to any activation function. Through extensive experiments, we find several important phenomena. Firstly, different activation functions (i.e., ReLU and SiLU) exhibit comparable performance but opposite training-time sparsity trends. The activation ratio (i.e., $1-\mathrm{sparsity\ ratio}$) evolves as a convergent increasing power-law and decreasing logspace power-law with the amount of training data for SiLU-activated and ReLU-activated LLMs, respectively. These demonstrate that ReLU is more efficient as the activation function than SiLU and can leverage more training data to improve activation sparsity. Secondly, the activation ratio linearly increases with the width-depth ratio below a certain bottleneck point, indicating the potential advantage of a deeper architecture at a fixed parameter scale. Finally, at similar width-depth ratios, we surprisingly find that the limit value of activation sparsity varies weakly with the parameter scale, i.e., the activation patterns within LLMs are insensitive to the parameter scale. These empirical laws towards LLMs with greater activation sparsity have important implications for making LLMs more efficient and interpretable. Materials related to this work (i.e., codes and checkpoints) are available at https://github.com/thunlp/SparsingLaw.
@@ 1 Introduction
Activation sparsity refers to the phenomenon where considerable elements within the output of a neural layer (typically activation functions, as shown in Figure 2) are zero or low values and thus contribute weakly to the final model output given a specific input. As a prevalent property of many language and vision modeling architectures (Li et al., 2022), activation sparsity has wide practical values, such as inference acceleration (Liu et al., 2023; Song et al., 2023; Xue et al., 2024; Song et al., 2024a), training acceleration (Zhang et al., 2024b), and LLM interpretation (Sajjad et al., 2022; Zhang et al., 2023). Generally, a model with a greater sparsity ratio (i.e., the ratio of inactivated elements) has more potential in these scenarios. However, this raises an underexplored problem: how to obtain an LLM with greater activation sparsity?
@@
Figure 1: A typical case of activation sparsity (with a sparsity ratio of 60%) in a gated feed-forward network of LLMs, where considerable elements weakly contribute to the outputs within the activation scores.
A simple solution is to design constraints at the model architecture level that force the model to have a predefined large sparsity ratio. For example, mixture-of-experts (MoE), the most popular architecture-constrained design, typically uses a token-level top-k parameter selection router to assign a fixed sparsity ratio for each token at each layer (Fedus et al., 2022; Zoph et al., 2022). However, these constraints often sacrifice model flexibility and performance. Recent works reveal the potential performance degradation caused by such inflexible sparsity assignment (Huang et al., 2024; Liu et al., 2024). Moreover, to inspect the impact of such constraints, we plot the PPL-activation (PPL denotes perplexity) Pareto curve ($\mathrm{activation\ ratio}=1-\mathrm{sparsity\ ratio}$) of MoE in Figure 2 and compare it with a vanilla decoder-only Transformer (Touvron et al., 2023) of the same parameter scale and amount of training data [^1] MoE models of different sparsity are obtained by tuning the number of activated experts, while for the vanilla setting, we adjust the CETT value proposed by Zhang et al. (2024a).. MoE has a significantly worse performance-sparsity trade-off. The best sparsity ratio is also hard to predefine, since a too-high or too-low sparsity ratio may lead to more severe performance degradation or substantial unnecessary computation, respectively.
To avoid negative impacts on flexibility and performance, we focus on the intrinsic activation sparsity within decoder-only Transformer-based LLMs in this paper, such as GPT (Brown et al., 2020) and LLaMA (Touvron et al., 2023). Different from the sparsity enforced through architectural constraints, intrinsic activation sparsity (Zhang et al., 2024a; Song et al., 2024a) arises self-adaptively during the pre-training stage, without unified or predefined sparsity ratios among different tokens or layers. We provide a comprehensive study on the quantitative scaling properties and influential factors of this intrinsic property. These findings can help us formulate a generalizable method to control and promote activation sparsity without harming flexibility and performance.
To rigorously study the influential factors of activation sparsity, we need a metric to precisely reflect the sparsity level of an LLM. Conventional works rely on prior experience to set a fixed global threshold for recognition of weakly-contributed neurons [^2] A neuron denotes a certain row or column within the parameter matrices., which is empirical, inflexible, and performance-unaware (see Section 2.2). We propose an improved metric, named PPL-$p\%$ sparsity, with three advantages: versatility across model architectures, performance-awareness, and the precise recognition of weakly-contributed neurons. For versatility, PPL-$p\%$ sparsity follows Zhang et al. (2024a) and recognizes the most weakly-contributed neurons, which are then inactivated, by introducing layer-wise adaptive thresholds and comparing the magnitudes of neuron outputs with the layer-specific threshold. As a magnitude-based metric, it is not bound to the activation functions that output exactly non-negative elements (e.g., ReLU) and is applicable to any LLMs with activation layers. Moreover, considering the trade-off between performance and sparsity (Song et al., 2024a), performance-awareness is required, indicating whether a metric can comprehensively reflect the sparsity under each desired performance level. To this end, PPL-$p\%$ sparsity explicitly incorporates PPL as a performance proxy, computed as the ratio of inactivated neurons when the output PPL raises just by $p\%$ after the layer-wise thresholds are applied, compared to the dense setting with all neurons activated. Finally, the precise recognition of weakly-contributed neurons indicates that our metric obtains a good trade-off between performance and sparsity ratio (see Section 4.1).
@@
Based on the above metric, we systematically study the correlation between the activation sparsity and influential factors, including the amount of training data, the activation function, the width-depth ratio (i.e., the ratio of the hidden dimension to the layer number), and the parameter scale. Through comprehensive experiments, we obtain the following observations:
@@
    There is an increasing power-law (SiLU-activated LLMs) or decreasing logspace power-law (ReLU-activated LLMs) relationship between the activation ratio and the amount of training data. Both laws are convergent with a certain limit sparsity ratio as the amount of data approaches infinity. Note that the increasing sparsity-data trend indicates that ReLU-activated LLMs are more efficient in improving activation sparsity with more data.
@@
    Given the same parameter scale, the sparsity obtained by ReLU-activated LLMs always surpasses that of SiLU-activated LLMs, while their performance is comparable.
@@
    Given the same parameter scale, the activation ratio linearly increases with the width-depth ratio under a bottleneck point (i.e., deeper models are sparser), above which the activation fluctuates around a fixed level. However, considering the performance issue, the best width-depth ratio should be just within a certain interval that ensures the best performance.
@@
    Given similar width-depth ratios, the limit of activation sparsity is weakly correlated to the scale of LLMs. On the other hand, the convergence speed to the limit is much faster in smaller models. We try to explain these phenomena in Section 4.4.
The above empirical laws can provide comprehensive instructional values for designing and pre-training an LLM with greater activation sparsity, which offers more significant potential in producing more efficient and interpretable LLMs. Moreover, our work enables the training-time prediction of the future sparsity ratio, and the evolving trend of activation sparsity with the amount of training data potentially provides a lens for the progress of neuron specialization.
@@ 2.1 Preliminaries of Activation Sparsity
Activation sparsity is a prevalent property existing in neuron networks with activation layers, indicating the existence of considerable parameters, which correspond to the activation outputs with zero or low values and thus have a limited impact on final network outputs given specific inputs. These weakly-contributed parameters are often defined as “inactivated parameters” and can be skipped in computation. Due to the activation layers in feed-forward networks (FFNs), mainstream LLMs also present remarkable activation sparsity (Li et al., 2022; Zhang et al., 2022; Song et al., 2024a).
Owing to activation sparsity, we can improve LLMs in many aspects, such as efficiency, interpretability, and robustness. For instance, recent works manage to exploit activation sparsity for inference acceleration, mainly by predicting the activation patterns, wisely allocating hardware resources, and saving redundant resources associated with inactivated parameters (Liu et al., 2023; Song et al., 2023; Xue et al., 2024)Zhang et al. (2024b) reveal the existence of activation sparsity throughout the majority of the LLM pre-training process, and then utilize this sparsity for pre-training acceleration. Besides acceleration, activation sparsity also helps improve the interpretability (Agarap, 2018; Dai et al., 2022; Cuadros et al., 2022; Sajjad et al., 2022) and robustness (Ahmad & Scheinkman, 2019; Muthukumar & Sulam, 2023) of LLMs, which are also important properties in producing reliable and well-performing LLMs.
@@ 2.2 Metrics of Activation Sparsity
While considerable works call for greater activation sparsity due to the significant merits brought by sparsity (e.g., more efficient computing and better model interpretability) (Mirzadeh et al., 2023; Song et al., 2024a; b), it is nontrivial to frame a satisfactory metric for measuring sparsity.
For the convenience of demonstrations, we formally introduce the following notations for the computation process of FFNs (also see Figure 2). By denoting the hidden dimension and intermediate dimension as $d_{h}$ and $d_{f}$ respectively, a gated FFN (a common FFN form adopted in most mainstream LLMs (Dauphin et al., 2017; Shazeer, 2020)) processes the inputs as follows:
@@
where $\mathbf{x}\in\mathbb{R}^{d_{h}}$, $\mathbf{s}\in\mathbb{R}^{d_{f}}$, $\sigma$, and $\odot$ denote the input hidden states, the activation scores, the activation function, and the element-wise multiplication, respectively. $\mathbf{W}^{gate},\mathbf{W}^{in}\in\mathbb{R}^{d_{f}\times d_{h}}$ and $\mathbf{W}^{out}\in\mathbb{R}^{d_{h}\times d_{f}}$ are learnable parameters. Next, we decompose the parameters of FFN along the dimension of $d_{f}$ into $d_{f}$ neurons. The output of the $i$-th neuron $n_{i}$ is calculated by
$s_{i}=\sigma(\mathbf{W}^{gate}_{i,:}\mathbf{x}),\quad n_{i}=\mathbf{W}^{out}_{:,i}[s_{i}\odot(\mathbf{W}^{in}_{i,:}\mathbf{x})],\quad\text{FFN}(\mathbf{x})=\sum_{i=1}^{d_{f}}n_{i},$ (2)
where $\mathbf{W}^{gate}_{i,:}$, $\mathbf{W}^{in}_{i,:}$, $\mathbf{W}^{out}_{:,i}$ are the $i$-th row of $\mathbf{W}^{in}$, the $i$-th row of $\mathbf{W}^{gate}$, and the $i$-th column of $\mathbf{W}^{out}$, respectively. The FFN outputs can be formalized as the sum of all neuron outputs.
Activation sparsity is measured by the ratio of inactivated neurons, namely $|\mathcal{D}|/d_{f}$, where $\mathcal{D}$ is the index set of inactivated neurons. However, different sparsity metrics can differ in determining whether a specific neuron is inactivated. A straightforward metric, naturally adopted in ReLU-activated models, regards those neurons with zero activation scores in Eq. (2) as inactivated, namely $\mathcal{D}=\{i|s_{i}=0\}$. To prune more non-zero weakly contributed activations, works by Kurtz et al. (2020) and Mirzadeh et al. (2023) introduce a positive threshold or bias. In this case, the set of inactivated neurons changes to $\mathcal{D}=\{i|s_{i}<\epsilon\}$, where $\epsilon>0$ is the threshold or bias.
The major drawback of this straightforward definition is the lack of versatility. Concretely, it is unsuitable for activation functions that have considerable non-negligible negative outputs, such as SiLU (Elfwing et al., 2018). In these cases, the straightforward metric can lose considerable negative neuron outputs and harm performance. A quick fix is to use the absolute value, $\mathcal{D}=\{i||s_{i}|<\epsilon\}$, but a global threshold across layers is hard to determine. To this end, Zhang et al. (2024a) adaptively searches the layer-wise thresholds by introducing the cumulative errors of tail truncation (CETT). Defined as the $L_{2}$ norm relative error caused by inactivated neurons, CETT is computed as:
@@
where $\|\cdot\|_{2}$ is the $L_{2}$ norm operator. Notably, as CETT increases monotonically with $\epsilon$, for each layer, we can use binary search to find a threshold $\epsilon$ that makes CETT just equal a predefined value.
Meeting the versatility demands, the CETT paradigm is still not friendly enough, as CETT does not directly reflect the model performance. In real-life deployment, we often need to utilize activation sparsity under a certain tolerance of performance degradation. Therefore, in this paper, we introduce a more performance-aware metric named PPL-$p\%$ sparsity, which explicitly reflects the sparsity ratio under a target PPL level by binary-searching an appropriate CETT value.
2.3 Scaling Properties of Activation Sparsity
@@
Despite the importance of activation sparsity, few works conduct comprehensive studies on how it scales with the increase of parameter scale and training data, as well as the impact of other influential factors within architecture designs. Speculating that activation sparsity comes from the training dynamic in the optimization process, Li et al. (2022) finds an increasing trend of sparsity with larger scale, depth, and width in T5 series (Raffel et al., 2020)Zhang et al. (2024a) dives into this problem from the aspect of activation functions. Song et al. (2024a) discovers that LLMs tend to be sparser on more formatted datasets such as codes and multiple choices. Other works have discussed the scaling properties of parameter-sparse models (Frantar et al., 2023), fine-grained MoE models (Krajewski et al., 2024), and sparse autoencoders (Gao et al., 2024).
To the best of our knowledge, we present the first comprehensive quantitative study on the scaling properties of activation sparsity and the impact of architecture factors. The work most similar to ours is done by Li et al. (2022), but they only conduct qualitative analyses and focus on the conventional encoder-decoder Transformer (e.g., T5) rather than decoder-only LLMs. Besides, they adopt the most straightforward metric for activation sparsity, limited to ReLU-activated models.
3 Methodology
3.1 Metric of Activation Sparsity
We propose PPL-$p\%$ sparsity, which improves the CETT metric (Zhang et al., 2024a) by directly reflecting the sparsity ratio at different requirements of PPL (as a proxy of performance). Firstly, given a pre-trained LLM, we introduce the dense setting with all its neurons activated (i.e., $|\mathcal{D}|=0$), whose outputs are theoretically the most accurate. Next, we conduct a binary search for an appropriate CETT value $\mathrm{CETT}_{k}$, which makes the average PPL increase just by $p\%$ compared to the dense setting. The search algorithm is described in Appendix C. Finally, based on $\mathrm{CETT}_{k}$, we can determine the activation thresholds for each layer and obtain the sparsity value, according to Eq. (3). Our experiments in Section 4.1 will demonstrate the rationality of this metric in achieving the greatest sparsity under the same output PPL.
3.2 Influential Factors
As we mainly study the effect of potentially influential factors on activation sparsity, it is necessary to specify how to compute the involved influential factors as follows:
  • Amount of training data: The number of tokens passed during the pre-training process. To obtain more comprehensive dynamics and more precise estimations of the limits, the number of training tokens is no less than 80 times the scale of non-embedding parameters.
  • Activation function: ReLU and SiLU are incorporated as the two most widely adopted activation functions with FFNs.
  • Parameter scale: The number of parameters included in a model excluding the embeddings. In this work, we incorporate 5 scales: 0.1B, 0.2B, 0.4B, 0.8B, and 1.2B.
  • Width-depth ratio: The ratio of the hidden dimension (i.e., $d_{h}$) to the number of hidden layers (denoted as $N$). We sample 9 values of width-depth ratio ranging from 14.2 to 597.3, covering both the ordinary and extreme conditions.
3.3 Setting of Scaling Analysis
Model settings
We adopt the same architecture as MiniCPM (Hu et al., 2024), which adopts Tensor Program (Yang et al., 2022) for training stability and shares the input and output embeddings.
Training settings

### v2

@@ Abstract
Activation sparsity denotes the existence of substantial weakly-contributed neurons within feed-forward networks of large language models (LLMs), providing wide potential benefits such as computation acceleration. However, existing works lack thorough quantitative studies on this useful property, in terms of both its measurement and influential factors. In this paper, we address three underexplored research questions: (1) How can activation sparsity be measured more accurately? (2) How is activation sparsity affected by the model architecture and training process? (3) How can we build a more sparsely activated and efficient LLM? Specifically, we develop a generalizable and performance-friendly metric, named CETT-PPL-1%, to measure activation sparsity. Based on CETT-PPL-1%, we quantitatively study the influence of various factors and observe several important phenomena, such as the convergent power-law relationship between sparsity and training data amount, the higher competence of ReLU activation than mainstream SiLU activation, the potential sparsity merit of a small width-depth ratio, and the scale insensitivity of activation sparsity. Finally, we provide implications for building sparse and effective LLMs, and demonstrate the reliability of our findings by training a 2.4B model with a sparsity ratio of 93.52%, showing 4.1$\times$ speedup compared with its dense version.
Machine Learning, ICML
luo-yq23@mails.tsinghua.edu.cn, scy22@mails.tsinghua.edu.cn
@@ 1 Introduction
Activation sparsity refers to the phenomenon that considerable elements within the output of activation layers (shown in Figure 1) are zero or low values, and thus the corresponding neurons [^1] A neuron denotes a certain row or column within the parameter matrices of feed-forward networks (FFN). contribute weakly to the final model output given a specific input. As a prevalent property of many language and vision modeling architectures (Li et al., 2022), activation sparsity has wide practical values, such as inference acceleration (Liu et al., 2023; Song et al., 2023), training acceleration (Zhang et al., 2024b), and LLM interpretation (Sajjad et al., 2022). For instance, Xue et al. (2024) achieves up to 27.8×27.8$\times$27.8 × inference acceleration on smartphones by utilizing activation sparsity, while Zhang et al. (2023) reveals and interprets the emergent modularity in LLMs by analyzing the specialization of sparsely-activated neurons.
@@
Figure 1: A typical case of activation sparsity (with a sparsity ratio of 60%) in a gated feed-forward network of LLMs.
Despite the extensive benefits of activation sparsity, there are few relevant comprehensive and quantitative studies. To fill this gap, we focus on three research questions:
  • (Q1) How can activation sparsity be measured more accurately?
  • (Q2) How is activation sparsity affected by the model architecture and training process?
  • (Q3) How can we build a more sparsely activated and efficient LLM?
Metric for Activation Sparsity
A “metric” for activation is for measuring the sparsity ratio (i.e., the ratio of weakly-contributed neurons) of an LLM given specific inputs. A good metric should be generalizable (i.e., applicable to different activation functions), and its impact on the model performance (after the neurons recognized as weakly-contributed ones are skipped during calculation) should be negligible. To achieve this, we need an accurate algorithm to identify weakly-contributed neurons in each layer. The more accurate this algorithm is, the less performance degradation a model will suffer under a specific sparsity ratio.
We experimentally demonstrate that CETT (Zhang et al., 2024a), a generalizable method adaptively setting layer-wise thresholds for activation values to identify weakly-contributed neurons (Section 2.2), achieves the best trade-off between performance and sparsity ratio among existing methods. To further ensure negligible performance degradation, we develop CETT-PPL-$1\%$ as a better sparsity metric. Specifically, this metric binary searches the CETT hyper-parameter (controlling the output error of FFNs in each layer after weakly-contributed neurons are skipped) under a PPL increase tolerance of 1% that causes nearly no harm to performance on downstream tasks. In the 0.8B ReLU-activated model (Figure 2), it reduces the activation ratio by 61% relatively compared with the common ReLU-oriented metric using a zero threshold.
Influential Factors of Activation Sparsity
@@
Using CETT-PPL-1%, we systematically study the correlation between activation sparsity and four influential factors, including the amount of pre-training data, the activation function, the width-depth ratio (i.e., the ratio of the hidden dimension to the layer number), and the parameter scale. Through comprehensive experiments on models scaled from 0.1B to 1.2B, we obtain the following observations:
@@
    There is an increasing power-law (mainstream SiLU-activated LLMs) or decreasing logspace power-law (ReLU-activated LLMs) relationship between the activation ratio ($1-\mathrm{sparsity\ ratio}$) and the amount of pre-training data. Both laws are convergent with a certain limit sparsity ratio as the amount of data approaches infinity. ReLU-activated LLMs can be more sparsely activated with more data. (Figure 4)
@@
    Given the same parameter scale, the sparsity ratio of ReLU-activated LLMs always surpasses that of SiLU-activated LLMs, with comparable performance on downstream tasks. (Figure 4 and Table 1)
@@
    Given the same parameter scale, the activation ratio linearly increases with the width-depth ratio under a bottleneck (i.e., deeper models are sparser), above which activation fluctuates around a fixed level. (Figure 6)
@@
    Given similar width-depth ratios, the limit of activation sparsity is weakly correlated to the scale of LLMs, while the convergence speed to the limit is much faster in smaller models. (Figure 8) We try to explain these phenomena in Section 5.3, indicating a similar neuron specialization pattern across models of distinct scales.
Approach to More Sparsely-activated LLMs
Based on the above findings, we answer the third question: To train an LLM with greater activation sparsity from scratch, it is better to adopt the ReLU activation, use a small width-depth ratio on the premise of stable training, and feed more training data so that the decreasing logspace power-law can help reduce the activation ratio.
To validate the generalizability of our findings, we train a larger 2.4B model. With ReLU activation, near 800B training tokens, and a width-depth ratio close to the 0.1B-1.2B experimental models (which is small enough and ensures training stability), the model undergoes a similar logspace power-law trend between activation and data, achieving a high limit sparsity ratio of 93.52% and 4.1$\times$ speedup ratio to the dense model. This value is also close to the limit sparsity ratios of 0.1B-1.2B models, consistent with the previously found scale insensitivity of the limit sparsity.
To sum up, our work provides a better practice of measuring activation sparsity and then comprehensively studies its influential factors and scaling properties. The empirical laws found above can provide instructional values for designing and pre-training an LLM with greater activation sparsity, which helps produce more efficient LLMs.
@@ 2.1 Preliminaries of Activation Sparsity
Activation sparsity is a prevalent property existing in neural networks with activation layers, indicating the existence of considerable neurons, that correspond to zero or low activation values, having a limited impact on final network outputs given specific inputs. Mainstream LLMs present remarkable activation sparsity (Li et al., 2022; Zhang et al., 2022; Song et al., 2025, 2024).
Activation sparsity can help improve LLMs in many aspects, such as efficiency and interpretability. Recent works manage to exploit activation sparsity for inference acceleration, mainly by saving the computation related to weakly-contributed parameters (Liu et al., 2023; Song et al., 2023; Xue et al., 2024)Zhang et al. (2024b) utilize the existence of activation sparsity throughout the majority of the LLM pre-training process to accelerate pre-training. Besides acceleration, activation sparsity also helps improve interpretability, which is important for reliable and well-performing LLMs. Explanation of the sparse neuron activation patterns is a mainstream paradigm of interpreting LLM behaviors (Dai et al., 2022; Sajjad et al., 2022; Bills et al., 2023; Gao et al., 2024; Lieberum et al., 2024).
Notably, mixture-of-experts (MoE) is a popular paradigm of activation sparsity, enforcing each token to activate a limited number of experts. As such constraints can sacrifice flexibility and performance (see Appendix B), we focus on activation sparsity in vanilla Transformers in this work.
Besides, activation sparsity also has fundamental differences from weight pruning, which realizes acceleration by removing certain parts of LLM parameters regardless of inputs. By contrast, activation sparsity adaptively determines important neurons for different inputs with no parameter permanently removed, causing considerably less performance degradation (see Appendix C).
@@ 2.2 Metrics of Activation Sparsity
Building a satisfactory metric for measuring sparsity is a nontrivial work. For the convenience of demonstrations, we formally introduce the following notations for the computation process of FFNs (also see Figure 1). With a hidden dimension $d_{h}$ and an intermediate dimension $d_{f}$, a gated FFN (the FFN paradigm in mainstream LLMs (Dauphin et al., 2017; Shazeer, 2020)) works as follows:
@@
where $\mathbf{x}\in\mathbb{R}^{d_{h}}$, $\mathbf{s}\in\mathbb{R}^{d_{f}}$, $\sigma$, and $\odot$ denote the input hidden states, the activation values, the activation function, and the element-wise multiplication, respectively. $\mathbf{W}^{gate},\mathbf{W}^{in}\in\mathbb{R}^{d_{f}\times d_{h}}$ and $\mathbf{W}^{out}\in\mathbb{R}^{d_{h}\times d_{f}}$ are learnable parameters. Next, we decompose the parameters of FFN along the dimension of $d_{f}$ into $d_{f}$ neurons. The output of the $i$-th neuron $n_{i}$ is calculated by
$s_{i}=\sigma(\mathbf{W}^{gate}_{i,:}\mathbf{x}),\quad n_{i}$ $=\mathbf{W}^{out}_{:,i}[s_{i}\odot(\mathbf{W}^{in}_{i,:}\mathbf{x})],$ (2)
$\text{FFN}(\mathbf{x})$ $=\sum_{i=1}^{d_{f}}n_{i},$
where $\mathbf{W}^{gate}_{i,:}$, $\mathbf{W}^{in}_{i,:}$, $\mathbf{W}^{out}_{:,i}$ are the $i$-th row of $\mathbf{W}^{in}$, the $i$-th row of $\mathbf{W}^{gate}$, and the $i$-th column of $\mathbf{W}^{out}$, respectively. The FFN outputs equals the sum of all neuron outputs.
Activation sparsity is measured by the ratio of weakly-contributed neurons, namely $|\mathcal{D}|/d_{f}$, where $\mathcal{D}$ is the index set of weakly-contributed neurons. Different metrics mainly differ in determining whether a specific neuron contributes weakly. A straightforward metric, naturally adopted in ReLU-activated models, regards neurons with zero activation values as weakly-contributed, namely $\mathcal{D}=\{i|s_{i}=0\}$. To find non-zero weakly contributed activations, Kurtz et al. (2020) and Mirzadeh et al. (2023) introduce a positive threshold or bias, i.e., $\mathcal{D}=\{i|s_{i}<\epsilon\},\ \epsilon>0$.
The major drawback of this straightforward definition is the lack of generalizability. Concretely, it is unsuitable for activation functions with considerable non-negligible negative outputs, such as SiLU (Elfwing et al., 2018). In these cases, the straightforward metric can lose considerable negative neuron outputs and harm performance. A quick fix is to use the absolute value, $\mathcal{D}=\{i||s_{i}|<\epsilon\}$, but a global threshold across layers is hard to determine. Zhang et al. (2024a) adaptively searches the layer-wise thresholds by introducing the cumulative errors of tail truncation (CETT). Defined as the $L_{2}$ norm relative error caused by skipping weakly-contributed neurons, CETT is computed as:
@@
where $\|\cdot\|_{2}$ is the $L_{2}$ norm operator. As CETT increases monotonically with $\epsilon$, for each layer, we can use binary search to find the threshold $\epsilon$ to meet the CETT value.
In this work, we experimentally demonstrate that CETT, when tuned to just make the validation PPL increase by 1%, is a metric with great generalizability and the least performance degradation, even at a high sparsity ratio.
2.3 Influential Factors of Activation Sparsity
@@
Despite the importance of activation sparsity, few works conduct comprehensive and quantitative studies on how it is affected by the model architecture and training process. Speculating that activation sparsity comes from the training dynamic in the optimization process, Li et al. (2022) finds an increasing trend of sparsity with larger scale, depth, and width in T5 series (Raffel et al., 2020)Zhang et al. (2024a) dives into this problem from the aspect of activation functions. Song et al. (2025) discovers that LLMs tend to be sparser on more formatted datasets such as codes and multiple choices. Other works have discussed the scaling properties of parameter-sparse models (Frantar et al., 2023), fine-grained MoE models (Krajewski et al., 2024), and sparse autoencoders (Gao et al., 2024). To the best of our knowledge, we present the first comprehensive quantitative study on the accurate measurement, influential factors, and training practice of activation sparsity for LLMs.
[…]
Figure 3: The Pareto curves between downstream task performance and activation sparsity using different methods to recognize weakly-contributed neurons. C.R. refers to commonsense reasoning and R.C. refers to reading comprehension.
3 Quantitative Study Settings
To address the three issues in the Introduction section, we conduct extensive experiments, training, evaluating, and analyzing the models ranging from 0.1B to 1.2B. We introduce our experimental settings in detail as follows.
Model settings We adopt the architecture of $\mu P$ Transformer (Hu et al., 2024), which combines the mainstream LLaMA (Touvron et al., 2023) with gated FFN and $\mu P$ parametrization (Yang et al., 2022) for training stability.